XYZ Electronics is a mid-sized electronics manufacturer located in Germany. The company president is Mark Johnson, who inherited the company. When it was founded, over 70 years ago, the company originally repaired radios and other household appliances. Over the years the company expanded into manufacturing, and is now a reputable manufacturer of various electronic items. Sam Smith, a recent MBA graduate, has been hired by the company’s finance department.

One of the major revenue-producing items manufactured by XYZ is a personal digital assistant (PDA). XYZ currently has one PDA model on the market, and sales have been excellent. The PDA is a unique item in that it comes in a variety of tropical colours and is pre-programmed to play Billy Bragg music. However, as with any electronic item, technology changes rapidly, and the current PDA has limited features in comparison with newer models. XYZ spent €750,000 to develop a prototype for a new PDA that has all the features of the existing PDA but adds new features such as cell-phone capability. The company has spent a further €200,000 for a marketing study to determine the expected sales figures for the new PDA.

XYZ can manufacture the new PDA for €155 each in variable costs. Fixed costs for the operation are estimated to be €4.7 million per year. The estimated sales volumes are 74,000, 95,000, 125,000, 105,000 and 80,000 per year for the next five years, respectively. The unit price of the new PDA will be €360. The necessary equipment can be purchased for €21.5 million, and will be depreciated using the 20 per cent reducing-balance method. It is believed the value of the equipment in five years will be €4.1 million.

As previously stated, XYZ currently manufactures a PDA. Production of the existing model is expected to be terminated in two years. If XYZ does not introduce the new PDA, sales will be 80,000 units and 60,000 units for the next two years, respectively. The price of the existing PDA is €290 per unit, with variable costs of €120 each and fixed costs of €1,800,000 per year. If XYZ does introduce the new PDA, sales of the existing PDA will fall by 15,000 units per year, and the price of the existing units will have to be lowered to €255 each. Net working capital for the PDAs will be 20 per cent of sales, and will occur with the timing of the cash flows for the year: for example, there is no initial outlay for NWC, but changes in NWC will first occur in year 1 with the first year’s sales. XYZ has a 35 per cent corporate tax rate and a 12 per cent required return.

Mark has asked Sam to prepare a report that answers the following questions.

Questions

a) What is the payback period of the project?

b) What is the profitability index of the project?

c) What is the IRR of the project?

d) What is the NPV of the project?
Answer #1
0 1 2 3 4 5
Sales in units of new PDA 74000 95000 125000 105000 80000
Sales revenue ($360) $   2,66,40,000 $3,42,00,000 $ 4,50,00,000 $3,78,00,000 $ 2,88,00,000
Variable cost ($155) $   1,14,70,000 $1,47,25,000 $ 1,93,75,000 $1,62,75,000 $ 1,24,00,000
Fixed costs (other than depreciation) $      47,00,000 $   47,00,000 $     47,00,000 $   47,00,000 $     47,00,000
Total Depn Book Value
Depreciation expense $      43,00,000 $   34,40,000 $     27,52,000 $   22,01,600 $     17,61,280 $     1,44,54,880 $      70,45,120
EBIT (New PDA) $      61,70,000 $1,13,35,000 $ 1,81,73,000 $1,46,23,400 $     99,38,720
Contribution lost on existing PDA:
Sales in units without new phone 80000 60000
Contribution magin at $290-$120 = $170 $   1,36,00,000 $1,02,00,000
Sales in units with new PDA 65000 45000
Contribution margin at $255-$120 = $135 $      97,50,000 $   67,50,000
Loss in contribution with new PDA $      38,50,000 $   34,50,000
Incremental EBIT $   1,00,20,000 $1,47,85,000 $ 1,81,73,000 $1,46,23,400 $     99,38,720
Tax at 35% $      35,07,000 $   51,74,750 $     63,60,550 $   51,18,190 $     34,78,552
NOPAT $      65,13,000 $   96,10,250 $ 1,18,12,450 $   95,05,210 $     64,60,168
Add: Depreciation $      43,00,000 $   34,40,000 $     27,52,000 $   22,01,600 $     17,61,280
OCF $   1,08,13,000 $1,30,50,250 $ 1,45,64,450 $1,17,06,810 $     82,21,448
Capital expenditure $     2,15,00,000
Change in NWC $   53,28,000.0 $   15,12,000 $     21,60,000 $ -14,40,000 $   -18,00,000 $        57,60,000
Release of NWC $    57,60,000
After tax salvage value = [4100000+(7045120-4100000)*35%] $     51,30,792
After tax annual cash flows $    -2,15,00,000 $      54,85,000 $1,15,38,250 $ 1,24,04,450 $1,31,46,810 $ 2,09,12,240
1) PAYBACK PERIOD:
Cumulative after tax cash flows $    -2,15,00,000 $ -1,60,15,000 $ -44,76,750 $     79,27,700 $2,10,74,510 $ 4,19,86,750
Payback period = 2+4476750/12404450 = 2.36 Years
2) PVIF at 12% [PVIF = 1/1.12^n] 1 0.89286 0.79719 0.71178 0.63552 0.56743
PV at 12% $      48,97,321 $   91,98,222 $     88,29,242 $   83,55,035 $ 1,18,66,167 $     4,31,45,988
PI = PV of cash inflows/Initial investment = 43145988/21500000 = 2.01
3) IRR:
PVIF at 41% [PVIF = 1/1.41^t] 1 0.70922 0.50299 0.35673 0.25300 0.17943
PV at 41% $    -2,15,00,000 $      38,90,071 $   58,03,657 $     44,25,070 $   33,26,166 $     37,52,364 $         -3,02,672
PVIF at 40% 1 0.71429 0.51020 0.36443 0.26031 0.18593
PV at 40% $    -2,15,00,000 $      39,17,857 $   58,86,862 $     45,20,572 $   34,22,223 $     38,88,305 $           1,35,820
IRR = 40%+1%*135820/(135820+302672) = 40.31%
4) NPV:
= PV of cash inflows-Initial investment = 43145988-21500000 $     2,16,45,988